OpenStudy (anonymous):

the diagonals of a rectangle STEP meet at K if SK=2x+6 and PK=36 find x?

OpenStudy (anonymous):

by the property of diagonal of rectangle, the pieces created when the diagonals intersect are congruent. therefore, SK=PK 2x+6=36 2x=36-6 2x=30 x=15